i need to list the vertex, focus, and directrix of this parabola
(x+2) = -8(y-3)^2

Answer by lwsshak3(11628)

i need to list the vertex, focus, and directrix of this parabola
(x+2) = -8(y-3)^2
**
Standard form of an equation for a parabola which opens leftwards: (y-k)^2=-4p(x-h), (h,k) being the (x,y) coordinates of the vertex.
..
(x+2) = -8(y-3)^2
rewrite equation
(y-3)^2=-(1/8)(x+2)
vertex:(-2,3)
Axis of symmetry: y=3
4p=1/8
p=1/32
focus:(-2-(1/32,3)=[-(65/32),3] (p units from vertex on the axis of symmetry)
directrix: x=-63/32 (a vertical line p units from vertex on the axis of symmetry)